OpenStudy (anonymous):

x + y = 600 .60x + .30y = .40(600)

OpenStudy (lilg132):

we use elimination in this y = 600 - x 0.60x + 0.30(600 - x) = 0.40(600) 0.60x + 180 - 0.30x = 240 0.60x - 0.30x = 240 - 180 0.30x = 60 x = 60/0.30 x = 200

OpenStudy (lilg132):

now we know x we can work out y x + y = 600 200 + y = 600 y = 600 - 200 y = 400 Dont forget to click best response

OpenStudy (anonymous):

30(i1-i2)-15i2=0 30i1+15i2=24

OpenStudy (lilg132):

post it again in another question